# Hardy–Sobolev interpolation inequalities

### v1

@@ 1. Introduction
$h[u]^{\theta}\|u\|_{2}^{2(1-\theta)}\geq C\|(-\Delta)^{q/2}u\|_{2}^{2},\quad\forall u\in H^{1}(\mathbb{R}^{d}),$ (3)
which was proved by Frank [6, Theorem 1.2]. Both of (2) and (3) have been extended to the fractional Laplacian in [6], motivated by applications in the asymptotic behavior of large Coulomb systems [19] and the stability of relativistic matter [8, 6].
@@
which was proved by Palatucci–Pisante [16, Theorem 1], using subtle weighted $L^{p}$-estimates for Riesz potentials in [18] and Calderón-Zygmund type techniques in the spirit of the Fefferman–Phong argument [5]. The bound (6) is helpful to obtain the compactness of minimizing sequences of the critical Sobolev inequality; see [16, Theorem 3] for details. In contrast, our inequality (4) is stronger than (5) and it only holds for the special case $d=1/\theta=3$.
@@
with a given constant $\eta>0$, were derived. In particular, according to [20, Eq. (4.6)], the only regular solutions (i.e., belonging to $L^{2^{*}}$), up to rescaling, are of the form $(d(d-2)\eta^{2})^{(d-2)/4}u_{\eta}$ with $u_{\eta}$ given in (10).
@@ Acknowledgements.
Part of the work has been done when CD was a visiting reseacher at the Institut des Hautes Études and she would like to thank Laure Saint-Raymond for her support and hospitality. We are grateful to Rupert Frank for helpful suggestions, leading to our consideration of Lorentz norms in Theorem 2. We also thank Hajer Bahouri, Morris Brooks, Albert Cohen, Patrick Gérard, Kihyun Kim, Mathieu Lewin, Hoai-Minh Nguyen, Nikita Simonov, Thomas Sørensen, Jakob Stern, Hanne van den Bosch and Jean van Schaftingen for interesting and inspiring discussions. We acknowledge the support from the Deutsche Forschungsgemeinschaft through the DFG project Nr. 426365943. CD also acknowledges the support from the Jean-Paul Gimon Fund and from the Erasmus+ programme.
@@ 2. Radial case
with $r=2^{*}=2d/(d-2)$, $\gamma=-1/r$. Actually it is a limiting case as $01/r+\gamma/n=$ 0 in dimension $n=1$, which does not seem available from the literature (see [13, Theorem 3.1] and [14, Theorem 2.2] for recent results in the limiting case).
@@
The optimal constant of the one-dimensional inequality (17) was already obtained by Nagy in 1941 [15], with $2^{*}=2d/(d-2)$ replaced by a general positive power. The existence and uniqueness of optmizers of the analogue of (17) in higher dimensions are also well-known; we refer to the classical works of Weinstein [23] and Kwong [11] for instance. The uniqueness of optimizers of (17) can be translated straightforwardly to the classification of optmizers of (9) as stated in Theorem 3 (i); we refer to [20, Eq. (4.6)] for a similar analysis.
@@
To have the equality in (11), we need to ensure all equalities in (2) and (23). The bound (23) contains two inequalities where the first equality occurs if there exists a constant $c>0$ such that
@@
while the second equality occurs if $|f|^{2}$ is monotone increasing on $(0,1)$ and monotone decreasing on $(1,\infty)$. Thus, we have all equalities in (23) if and only if
@@ 3.1. Proof of Theorem 1
and $A-B\geq\theta A$, we conclude from (6) (see [16, Theorem 1]) that
@@ 3.2. Proof of Theorem 2
From [16, Theorem 1], see also [17, Eq. (1.4)], we have

### v2

@@ 1. Introduction
$h[u]^{\theta}\|u\|_{2}^{2(1-\theta)}\geq C\|(-\Delta)^{s/2}u\|_{2}^{2},\quad\forall u\in H^{1}(\mathbb{R}^{d}),$ (3)
for $s\in(0,1)$ and $\theta=\theta(s)$, which was proved by Frank [6, Theorem 1.2]. Both of (2) and (3) have been extended to the fractional Laplacian in [6], motivated by applications in the asymptotic behavior of large Coulomb systems [20] and the stability of relativistic matter [8, 6].
@@
which was proved by Palatucci–Pisante [17, Theorem 1], using subtle weighted $L^{p}$-estimates for Riesz potentials in [19] and Calderón-Zygmund type techniques in the spirit of the Fefferman–Phong argument [5]. The bound (6) is helpful to obtain the compactness of minimizing sequences of the critical Sobolev inequality; see [17, Theorem 3] for details. In contrast, our inequality (4) is stronger than (5) and it only holds for the special case $d=1/\theta=3$.
@@
with a given constant $\eta>0$, were derived. In particular, according to [21, Eq. (4.6)], the only regular solutions (i.e., belonging to $L^{2^{*}}$), up to rescaling, are of the form $(d(d-2)\eta^{2})^{(d-2)/4}u_{\eta}$ with $u_{\eta}$ given in (10).
@@ Acknowledgements.
Part of the work has been done when CD was a visiting reseacher at the Institut des Hautes Études and she would like to thank Laure Saint-Raymond for her support and hospitality. We are grateful to Rupert L. Frank for helpful suggestions, leading to our consideration of Lorentz norms in Theorem 2. We also thank Hajer Bahouri, Morris Brooks, Albert Cohen, Patrick Gérard, Kihyun Kim, Mathieu Lewin, Hoai-Minh Nguyen, Julien Sabin, Nikita Simonov, Thomas Sørensen, Jakob Stern, Hanne van den Bosch and Jean van Schaftingen for interesting and inspiring discussions. We acknowledge the support from the Deutsche Forschungsgemeinschaft through the DFG project Nr. 426365943. CD also acknowledges the support from the Jean-Paul Gimon Fund and from the Erasmus+ programme.
@@ 2. Radial case
with $r=2^{*}=2d/(d-2)$, $\gamma=-1/r$. Actually it is a limiting case as $01/r+\gamma/n=$ 0 in dimension $n=1$, which does not seem available from the literature (see [13, Theorem 3.1], [14, Theorem 1.2] and [15, Theorem 2.2] for recent results in the limiting case).
@@
The optimal constant of the one-dimensional inequality (17) was already obtained by Nagy in 1941 [16], with $2^{*}=2d/(d-2)$ replaced by a general positive power. The existence and uniqueness of optmizers of the analogue of (17) in higher dimensions are also well-known; we refer to the classical works of Weinstein [24] and Kwong [11] for instance. The uniqueness of optimizers of (17) can be translated straightforwardly to the classification of optmizers of (9) as stated in Theorem 3 (i); we refer to [21, Eq. (4.6)] for a similar analysis.
@@
To have the equality in (11), we need to ensure all equalities in (2) and (LABEL:eq:cs-inf-f). The bound (LABEL:eq:cs-inf-f) contains two inequalities where the first equality occurs if there exists a constant $c>0$ such that
@@
while the second equality occurs if $|f|^{2}$ is monotone increasing on $(0,1)$ and monotone decreasing on $(1,\infty)$. Thus, we have all equalities in (LABEL:eq:cs-inf-f) if and only if
@@ 3.1. Proof of Theorem 1
and $A-B\geq\theta A$, we conclude from (6) (see [17, Theorem 1]) that
@@ 3.2. Proof of Theorem 2
From [17, Theorem 1], see also [18, Eq. (1.4)], we have